The WireOne Company manufactures high-quality coated electrical wire in two departments, Weaving and Coating. Materials are introduced at various points during work in the Weaving Department. After the weaving is completed, the materials are transferred into the Coating Department, where specialty plastic coating is applied.

Selected data relating to the Weaving Department during May are given below:

  Production data:
    Kilograms in process, May 1(materials 100%
          complete; conversion 80% complete)
127,000
    Kilograms started into production during May 462,000
    Kilograms completed and transferred to Coating ?
    Kilograms in process, May 31 (materials 65%
          complete; conversion 30% complete)
74,000
  Cost data:
    Work in process inventory, May 1:
        Materials cost $ 118,110
        Conversion cost $ 87,630
    Cost added during May:
        Materials cost $ 861,684
        Conversion cost $ 304,526

The company uses the weighted-average method.

Required:

1. Compute the equivalent units of production.

2. Compute the costs per equivalent unit for May. (Round your answers to 2 decimal places.)

3. Determine the cost of ending work in process inventory and of the units transferred to the Coating Department. (Round intermediate calculations to 2 decimal places, and final answers to the nearest whole dollar.)

4. Prepare a cost reconciliation between the costs determined in (3) above and the cost of beginning inventory and costs added during the period. (Round intermediate calculations to 2 decimal places, and final answers to the nearest whole dollar.)
Answer #1

1. Equivalent units of production = Units transferred out + Ending WIP x Completion
Units completed and transferred = 127000+462000-74000 = 515000 kilograms

Equivalent units of Production
Material = 515000 + 74000 x 65% = 563100 kilograms
Conversion = 515000 + 74000 x 30% = 537200 kilograms

2. Cost per equivalent unit = (Beginning Cost + Cost added) / Equivalent units
Material = ($118110+861684)/563100 = $1.74 per kilogram
Conversion = ($87630+304526)/537200 = $0.73 per kilogram

3.
Cost of Ending Work in process inventory = $83694 + $16206 = $99900
Material = 74000 x 65% x $1.74 = $83694
Conversion = 74000 x 30% x $0.73 = $16206

Cost of Units transferred = $896100 + $375950 = $1272050
Material = 515000 x $1.74 = $896100
Conversion = 515000 x $0.73 = $375950

4.

Reconciliation of Costs
Costs to be accounted for
Beginning WIP $       9,79,794
Cost added $       3,92,156
Total Costs to be accounted for $    13,71,950
Costs accounted for
Ending WIP $          99,900
Units transferred $     12,72,050
Total Costs accounted for $    13,71,950